Iodine-131, t_1/2 = 8.0 days, is used in diagnosis
Answer #1

Radio active decay is a first order reaction.

For first order recation,

half life t1/2 = 0.693 /k where k is rate constant

k = 0.693/ t1/2    ------ Eq (1)

k = 1/t ln { [A]o/[A]t}    ----- Eq (2)

From Eqs (1) and (2),

0.693/ t1/2 = (1/t) ln {[A]o/ [A]t} ------Eq (3)

Given that

half life of I-131, t1/2= 8.0 days

time t = ?

Initial activity of I-131 [A]o = 9.95 x 1018

Final activity of I-131   [A]t = 6.22 x 1017

Substitute all the values in Eq (3),

0.693/ t1/2 = (1/t) ln {[A]o/ [A]t}

[0.693/8] hrs = (1/t )  ln {9.95 x 1018/ 6.22 x 1017}

Then,

t = [ 8/0.693] x ln {9.95 x 1018/ 6.22 x 1017}

= 32 days

t = 32 days

Therefore,

required time = 32 days